solve each equation.
3(x + 2) = 5(x - 6)

Answer by TimothyLamb(4379)

3x + 6 = 5x - 30
5x - 3x = 6 + 30
2x = 36
x = 36 / 2
x = 18
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php